OpenStudy (landon34):

Which of the following is a solution to the equation x-5/7=-3/7. need some help.

OpenStudy (benlindquist):

awww, bruh

OpenStudy (kendricklamar2014):

Is this your question: \[x - \frac{ 5 }{ 7 } = - \frac{ 3 }{ 7 }\]

OpenStudy (landon34):

yes

OpenStudy (landon34):

i got stuck on a question when I was doing homeschool work

OpenStudy (kendricklamar2014):

Ok, here is how you solve this: Step 1: Simplify: \[x + \frac{ -5 }{ 7 } = \frac{ -3 }{ 7 }\] Step 2: Add 5/7 to both sides: \[x + \frac{ -5 }{ 7 } + \frac{ 5 }{ 7 } = \frac{ -3 }{ 7 } + \frac{ 5 }{ 7 }\] \[x = \frac{ 2 }{ 7 }\]

OpenStudy (landon34):

ok thanks. thats all I need.

OpenStudy (kendricklamar2014):

Your Welcome :)